OpenStudy (anonymous):

find the values of b and c which will make f(x) continuous for all real numbers if f(x) = x+1 abs(x-2) <1 x^2+bx+c abs(x-2) >=1

OpenStudy (anonymous):

its the absolute values that are throwing me off. it's in the left braces and the second one is greater than or equal to, sorry for any confusion.

OpenStudy (anonymous):

first off what does \[|x-2|<1\] really mean? it means \[-1<x-2<1\] i.e. \[1<x<3\] so we are only concerned with the points x = 1 and x = 3 where the definition of the function changes

OpenStudy (anonymous):

so here is what you do. replace x by 1 in both formulas and set them equal. you get \[1+1=1+b+c\]

OpenStudy (anonymous):

so you know \[b+c=1\]

OpenStudy (anonymous):

To be continuous, the two pieces must equal the same value when one piece switches to another (i.e. - x={1,3}). Thus:\[(1)+1=(1)^2+b(1)+c \text{ and }(3)+1=(3)^2+b(3)+c\]\[1=b+c\text{ and } -5=3b+c\]The solution of this system of equations is:\[b=-3 \text{ and } c=4\]

OpenStudy (anonymous):

now replace x by 3 in both formulas. you get \[3+1=9+3a+c\] so \[3a+c=-5\]

OpenStudy (anonymous):

damn i used a instead of b sorry i will rewrite it

OpenStudy (anonymous):

It's done already, dude.

OpenStudy (anonymous):

that makes so much sense, thank you!

OpenStudy (anonymous):

you get \[3+1=9+3b+c\] so \[3b+c=-5\] now solve the system \[b+c=1\] \[3b+c=-5\] etc